OpenStudy (anonymous):

Evaluate 2x-3y+z, if x=2, y=-2, and z=5 A. 3 B. -3 C. 15 D. 0

OpenStudy (anonymous):

its c

OpenStudy (zpupster):

just plug in 2(2)-3(-2)+5=4+6+5=